OpenStudy (anonymous):

Adding and Subtracting polynomials: (3a +2b-7c)+ (6b-4a+9c) + (-7c-3a-2b) I think I need it right but could someone give me the answer so I can double check??

OpenStudy (anonymous):

did*

jimthompson5910 (jim_thompson5910):

what answer did you get?

OpenStudy (anonymous):

its in my notebook at school. I'll have to check it in the morning!!

jimthompson5910 (jim_thompson5910):

How are you going to check if it's all the way over there? Anyways, if you add the 'a' terms, you get 3a + (-4a) + (-3a) = ???

OpenStudy (anonymous):

I have school in the morning.. first hour is studyhall! I think I did get something like that with the 3a + (-4a) thank you!

jimthompson5910 (jim_thompson5910):

what does 3a + (-4a) turn into when you add?

OpenStudy (anonymous):

-1a right?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

then we add on (-3a) so -1a + (-3a) = -4a

jimthompson5910 (jim_thompson5910):

therefore, we know this so far 3a + (-4a) + (-3a) = -4a

jimthompson5910 (jim_thompson5910):

now onto the b terms what is 2b+6b + (-2b) equal to?

OpenStudy (anonymous):

6b ?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

and finally, the c terms what do you get when you add up all the c terms?

OpenStudy (anonymous):

-5c ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

all together, we know that (3a +2b-7c)+ (6b-4a+9c) + (-7c-3a-2b) simplifies to -4a+6b-5c

OpenStudy (anonymous):

okay, that makes sense. So our final answer is -4a+6b-5c

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

Okay! thank you!

jimthompson5910 (jim_thompson5910):

you're welcome